OpenStudy (anonymous):

The Sweet Candy Shop buys 600 pounds of chocolate covered strawberries at $5.59 per pound. If a 10% spoilage rate is anticipated, at what price per pound should the strawberries be sold in order to achieve a 60% markup based on cost?

OpenStudy (anonymous):

0.9 * 600 * x = 600 * 5.59 * 1.6 x = (600 * 5.59 * 1.6) / 0.9 * 600 = $9.94 per pound

OpenStudy (anonymous):

thanks